OpenStudy (kj4uts):

Suppose a professor announces the following grading policy to her class: there will be four 1 hour exams and a final. In computing the course average the lowest grade on the hour exams may be replaced by the grade on the final exam if the grade on the final exam is higher. Assume all exams have scores between 0 and 100, all tests are equally weighted, and the course average is the usual average of all exams and the final. Also assume the course grade is based on the course average and that the standard scale applies. 90-100 = A 80-89 = B 70-79 = C 60-69 = D 0-59 = F

OpenStudy (kj4uts):

Also assume all tests and the final are equally weighted. If the student’s scores on the hour exams are 72, 85, 82 and 78, and the student gets 76 on the final, determine the course final grade for the student.

OpenStudy (kj4uts):

I got 80.25 because 76(final)+85+82+78/4=80.25 I left out the 72 because it is supposed to be replaced by the final but maybe I am not reading the question right.

OpenStudy (kj4uts):

If I included 72+85+82+78+76=393/5=78.6 or C

OpenStudy (anonymous):

The 72's not dropped, it's changed to the grade on the final. (76 + 85 + 82 + 78 + 76)/5 = 79.4

OpenStudy (anonymous):

the LOWEST grade on the hour exams may be REPLACED BY the grade on the FINAL exam if the grade on the final exam is higher

OpenStudy (kj4uts):

oh I see so the grade would be c then

OpenStudy (anonymous):

Yes

OpenStudy (kj4uts):

Thanks!